# Diffusion-Free Dynamics in Rotating Spherical Shell Convection Driven By Internal Heating and Cooling

### v1

@@ 3 Results
Figure 3: Snapshots of the perturbation temperature for six experiments. The experiments shown in the top row are in the rotationally-uninfluenced regime, those in the middle row are in the rotationally-influenced regime, and those in the bottom row are in the rapidly rotating regime. In each panel, the inner surface corresponds to $r=4.2$ and the outer surface corresponds to $r=4.8$.
@@ 3.1 Simulation results
Figure 2a shows the shell-averaged temperature difference $\Delta T$ between the inner and outer radii of the CZ. For $Ro_{\text{cv,F}}<1$, $\Delta T$ displays a strong dependence on $Ro_{\text{cv,F}}$, increasing with $\Omega_{0}$ as the rotational constraint acts to throttle convective heat transport. For each $Ro_{\text{cv,F}}$, the scatter between experiments with different $Ra_{\text{F}}$ is very small. In the rotationally-influenced regime, the dependence of $\Delta T$ on $Ro_{\text{cv,F}}$ ($\Delta T\propto\Omega_{0}^{\frac{4}{5}}\,{\sim}\,Ro_{\text{cv,F}}^{-\frac{4}{5}}$) is consistent with rotating mixing length theory (RMLT; see Section 3.2; D. J. Stevenson, 1979; A. J. Barker et al., 2014), which is a diffusion-free theory. For the cases categorised as rotationally-constrained with $Ro_{\text{cv,F}}<0.1$, $\Delta T$ depends more strongly on $Ro_{\text{cv,F}}$ than predicted by RMLT, while remaining independent of the diffusivity (for the values of $Ro_{\text{cv,F}}$ for which there are multiple experiments). In the rotationally-uninfluenced regime where $Ro_{\text{cv,F}}>1$, the rotational constraint is lost, causing $\Delta T$ to become independent of $Ro_{\text{cv,F}}$. Our simulations yield results that are consistent with this expectation, and thus by extension are consistent with diffusion-free behaviour.
@@
The azimuthal velocity averaged within $\pm 5^{\circ}$ of the equator and throughout the depth of the CZ, $U_{\text{eq}}$, is shown in Figure 2b. The dependence of $U_{\text{eq}}$ on $Ro_{\text{cv,F}}$ we obtain is consistent with previous work (T. Gastine et al., 2014). In the absence of a rotational constraint ($Ro_{\text{cv,F}}>1$), differential rotation does not develop ($U_{\text{eq}}\approx 0$). In the rotationally-influenced regime ($Ro_{\text{cv,F}}<1$; $Ro_{\omega}\gtrsim 1$), anti-Solar (retrograde) differential rotation develops, with an amplitude that increases as the rotation rate is increased. In this regime, the amplitude of the equatorial zonal flow is independent of the diffusivity. When rotation becomes dominant ($Ro_{\text{cv,F}}<0.1$; $Ro_{\omega}\lesssim 1$), the differential rotation switches direction from retrograde to prograde (Solar-like). In this regime, the data in general do not collapse onto a single curve for $U_{\text{eq}}$ vs. $Ro_{\text{cv,F}}$ indicating that, at least for the supercriticalities considered in the present work, the strength of $U_{\text{eq}}$ depends on the diffusivities. We note that there are only two values of $Ro_{\text{cv,F}}<0.1$ for which we ran a significant number of experiments with varying $Ra_{\text{F}}$ (5 experiments each for $Ro_{\text{cv,F}}=3.11\times 10^{-2}$ and $6.70\times 10^{-2}$). For the series with $Ro_{\text{cv,F}}=6.70\times 10^{-2}$, $U_{\text{eq}}$ does not vary significantly as the diffusivity is varied, but for the series with $Ro_{\text{cv,F}}=3.11\times 10^{-2}$ there is a diffusivity dependence.
@@
Figures 2c and d show the kinetic energy (KE) averaged over the volume of the CZ. Figure 2c shows $\text{KE}_{\text{fluc}}$, computed using the r.m.s. fluctuating velocity, and Figure 2d shows the ratio DRKE/KE, where KE is the total kinetic energy, and DRKE is the ‘differential rotation kinetic energy’, computed using the axially-symmetric component of the azimuthal velocity. The behaviour of $KE_{\text{fluc}}$ in Figure 2c tells a similar story to $U_{\text{eq}}$, namely that the kinematic properties of the flow display a weak dependence on the diffusivity for $Ro_{\text{cv,F}}>0.1$ (rotationally-uninfluenced or rotationally-influenced), but a stronger dependence once the influence of rotation becomes dominant. There are two additional features of Figure 2c,d that are worth emphasising. First, we note that in the rotationally-influenced regime, the dependence of $KE_{\text{fluc}}$ on $Ro_{\text{cv,F}}$ is consistent with the diffusion-free ‘Coriolis-inertial-Archimedean’ (CIA) scaling (e.g., T. Gastine et al., 2016; see Section 3.2). Second, we highlight that while $\text{KE}_{\text{fluc}}$ (Figure 2c) and the DRKE depend on the diffusivity in the rotationally-constrained regime, they do so in such a way that the ratio DRKE/KE is roughly diffusion-free (Figure 2d). In this regime, the DRKE dominates the total KE so that their ratio saturates at $\text{DRKE}/\text{KE}\approx 1$.
@@
In the rotationally-uninfluenced regime, convective plumes extend outwards in the radial direction at all latitudes and the effect of rotation is undetectable. In this regime, the temperature field is organised into large ‘warm’ and ‘cool’ patches that are dominated by spherical harmonic degree $\ell=1$ or $2$. The morphology in the rotationally-influenced cases with $Ro_{\text{cv,F}}=3.11\times 10^{-1}$ is similar, although the convection at higher latitudes now exhibits a weak alignment with the rotation axis and some low-latitude features are elongated in the azimuthal direction. However, the impact of rotation is hard to detect, consistent with $Ro_{\text{bulk}}>1$. Figure 4 shows an additional case in the rotationally-influenced regime with $Ro_{\text{cv,F}}=1.44\times 10^{-1}$ (i.e., a slightly faster rotation rate), which more clearly shows that the convection is aligned cylindrically, and zonally-elongated at lower latitudes. We highlight that the pair of experiments with $Ro_{\text{cv,F}}=3.11\times 10^{-1}$ are morphologically very similar, as are the pair with $Ro_{\text{cv,F}}=1.44$, which is consistent with the inference from Figure 2 that the diffusivity has a limited influence on the dynamics in the rotationally-uninfluenced and rotationally-influenced regimes.
The effect of rotation in the rotationally-constrained regime is clear. In mid-latitudes and the polar regions, convection is aligned cylindrically (i.e., with the rotation axis). Closer to the equator, large, modulated thermal Rossby wave-like features can be identified, that propagate in the prograde direction (with respect to the mean zonal flow). These waves are present in all experiments with $Ro_{\text{cv,F}}<0.1$ but in none with $Ro_{\text{cv,F}}\geq 1$. The transition to a regime of equatorial dynamics dominated by wave-like features occurs when the small scales become rotationally constrained ($Ro_{\omega}<1$). For the relatively turbulent cases shown in Figure 3 there is a strong polar vortex. This is not present in simulations that are more strongly dissipative (an example is given by the lower panel of Figure 4).
@@
Each of the scatter charts in this article use the marker shape to illustrate the occurrence of the SF, FSF, and FS configurations of differential rotation in parameter space (circles for SF, pentagons for FSF, and squares for FS). From Figure 1a–c, we identify that solar-like differential rotation (FS) is only obtained for the most diffusive experiments ($Ra_{\text{F}}\lesssim 10^{6}$), while more turbulent simulations with a strong rotational constraint exhibit FSF differential rotation. Figure 1d shows that the transition from a fast equator (either FS or FSF) to a slow equator occurs at the boundary between the rotationally-constrained and rotationally-influenced regimes, when $Ro_{\text{cv,F}}=0.1$, which corresponds to $Ro_{\omega}\approx 1$, consistent with previous work (T. Gastine et al., 2014). The fact that all of the FS or FSF cases have $Ro_{\text{cv,F}}<0.1$, and all of the SF cases have $Ro_{\text{cv,F}}\geq 0.1$, implies that the transition from a fast equator to a slow equator is diffusion-free (even if the zonal flow structure in the fast-equator cases themselves displays a dependence on the diffusivity). We believe that this transition is associated with the existence of thermal Rossby wave-like features at low latitudes, which are present in all simulations in the rotationally-constrained regime but absent otherwise. However, a complete analysis of the zonal momentum budget is left for future work.
@@
Figure 6: Scaling results for NuNu (panels a and b) and ReRe (panels c and d) obtained with our experiments. In panels a) and b), black and grey dashed lines correspond to the mixing length theory (MLT) and rotating mixing length theory (RMLT) scaling laws given by Equations 12 and 13, respectively, and the black solid lines connect selected series of experiments with constant $Ro_{\text{cv,F}}$ (see caption of Figure 1 a–c for further information). In panel c), the scaling for ReRe predicted by VAC balance (Equation 17) is shown as a black dashed line. Finally, in panel d), the black dashed line shows the inertial scaling of rotating convection (associated with CIA balance; given by Equation 16), and the grey dashed line shows the diffusion-free, non-rotating ‘ultimate’ scaling $Re\propto Ra^{1/2}$ (Equation 15).
@@ 3.2.1 Nusselt number scaling
for $Pr=1$. A central assumption in MLT is that the total dimensional heat transport is independent of the diffusivities, and thus the scaling above is diffusion-free. We note that a simple dimensional analysis for a non-rotating system yields the scaling above under this assumption [^2] A diffusion-free scaling for a non-dimensional number (e.g., $Nu$ or $Re$) does not mean that the scaling parameter (e.g., $Ra$) is diffusivity-independent. Instead, it is a scaling that is consistent with the dimensional quantities (e.g., the dimensional radial temperature gradient for $Nu$) being diffusivity-independent. For the specific case of the MLT $Nu$ scaling, $Nu$ includes a factor of $\kappa^{-1}$ if $T$ is re-dimensionalised, and so the right-hand-side must also contain a factor of $\kappa^{-1}$, which implies a scaling $Nu\propto Ra^{\frac{1}{2}}Pr^{\frac{1}{2}}$ (which is $Ra^{\frac{1}{2}}$ for $Pr=1$).. For a rotating system, the analogous theory to MLT is rotating mixing length theory (RMLT; D. J. Stevenson, 1979), which assumes that convection is dominated by modes with a growth rate well described by linear theory and amplitudes determined by a balance between their growth rate and the non-linear cascade rate. RMLT predicts that $Nu$ scales as
@@
As with Equation 12, this is a diffusion-free scaling. It scaling can also be obtained by combining the ‘diffusion-free heat transport assumption’ with a further assumption that the heat transport depends only on the supercriticality of the system $Ra/Ra_{\text{c}}$ (recall that $Ra_{\text{c}}\,{\sim}\,Ta^{\frac{3}{2}}$ for large $Ta$) (K. Julien et al., 2012; A. J. Barker et al., 2014). We note that $Ra\approx Ra_{\text{F}}/Nu$, which yields the scaling $Nu\propto\left(Ra_{\text{F}}Ta^{-2/3}\right)^{3/5}Pr^{-1/5}$. Using the definition of $Ro_{\text{cv,F}}$, this can be re-written $Nu\propto Ra_{\text{F}}^{1/3}Pr^{1/3}Ro_{\text{cv,F}}^{4/5}$. Therefore, with $Pr$ constant, RMLT implies
@@
(i.e., as in MLT). Previous studies of boundary-driven Boussinesq convection in a spherical shell have not yielded MLT-like scaling for the Nusselt number in the regime where rotation is unimportant. Instead, shallower scaling exponents were obtained, due to the throttling of convection by the conductive heat transport in the boundary layers (e.g., T. Gastine et al., 2015). Some studies have obtained RMLT-like scaling in the rapidly rotating regime, but only for extreme parameter values (e.g., T. Gastine et al., 2016, wherein experiments with $Ta\gtrsim 10^{12}$ and $RaTa^{-2/3}\simeq 10$ exhibit a Nusselt number scaling in agreement with RMLT).
@@
Inspection of Figure 6a reveals that our experiments fall into three groups distinguished by different $Nu$ vs. $RaTa^{-2/3}$ scaling behaviours. When $Ro_{\text{cv,F}}\geq 1$ (indicated by a green marker colour), the Nusselt number follows the MLT scaling prediction (Equation 12). These experiments are those that we have categorised as rotationally-uninfluenced. For, $0.1\leq Ro_{\text{cv,F}}<1$ (indicated by blue marker colours), i.e., experiments in the rotationally-influenced regime, a steeper scaling is obtained that is consistent with RMLT (Equation 13). The transition between the MLT and RMLT scalings occurs when $Ro_{\text{cv,F}}=1$, which roughly corresponds to the point where $Ro_{\text{bulk}}=1$. The agreement between the scaling obtained by the rotationally-influenced experiments and RMLT is striking, given that the effect of rotation on the convective morphology in this regime can be relatively weak (e.g., Figure 3). These results (as well as those presented in the next subsection) inform our definition of the rotationally-uninfluenced and rotationally-influenced regimes. Finally, in the rotationally-constrained regime for which $Ro_{\text{cv,F}}<0.1$ (orange marker colours), the Nusselt number scaling departs from the RMLT branch occupied by the rotationally-influenced cases. This departure from RMLT is manifest as an offset from the RMLT branch (reduced $Nu$ for a given $RaTa^{-2/3}$) that increases as $Ro_{\text{cv,F}}$ is reduced. For constant $Ro_{\text{cv,F}}$, experiments in this regime roughly follow the $Nu\propto Ra^{1/2}$ scaling noted above (Equation 14; for simulations with sufficient supercriticality, $RaTa^{-2/3}\gtrsim 5$). This indicates that the convective heat transport is still independent of the diffusivity (consistent with the pseudodimensional $\Delta T$ shown in Figure 2a), and that the offset from the rotationally-influenced RMLT branch captures a dependence of the convective heat transport on $Ro_{\text{cv,F}}$ (a diffusion-free parameter) that is not accounted for by RMLT. Hereafter, we refer to this scaling as ‘offset RMLT’. As noted previously, the transition from the rotationally-influenced regime to this new regime occurs when $Ro_{\omega}\lesssim 1$, which indicates that all spatial scales are now affected by rotation, motivating our choice to refer to this regime as rotationally-constrained.
@@
When the same data are plotted against $Ra$ (Figure 6b), the MLT scaling obtained by the rotationally-uninfluenced regime can be identified for experiments with $Ta=10^{5}$, which unambiguously follow the $Ra^{1/2}$ prediction. The $Ta=10^{6}$ series follows the RMLT scaling for moderate supercriticality, before transitioning towards the MLT scaling at higher supercriticality, and the experiments at $Ta=10^{7}$ are approaching the RMLT scaling. Finally, the offset RMLT scaling identified for the rotationally-constrained regime is manifest as a steepening of the $Nu(Ra)$ relation for fixed $Ta$ (but, again, note that series of experiments with constant $Ro_{\text{cv,F}}$, joined by black solid lines, follow the RMLT prediction given by Equation 14).
@@
A preliminary analysis (not shown) of the latitudinal dependence of the heat budget and Nusselt number (cf. T. Gastine & J. M. Aurnou, 2023) suggests that the transition to the offset RMLT scaling that characterises the rotationally-constrained regime may be related to the change in the dynamics that from a flow dominated by turbulent convection to a wave-dominated flow (we find wave-like features are always present when $Ro_{\text{cv,F}}<0.1$, but are absent when $Ro_{\text{cv,F}}\geq 0.1$). However, the offset scaling could alternatively arise due to interaction between the convection and the zonal flow, although this would have to be a ‘signed’ effect, since strong (anti-Solar) zonal flows are also present in the rotationally-influenced regime for which the standard RMLT scaling is recovered. We will explore this further in future work. We note that the offset scaling persists even for the rotationally-constrained experiments with the highest supercriticality. As a result, we do not believe that it arises because the rotationally-constrained simulations generally have lower supercriticality than those categorised as rotationally-influenced.
@@
Note. — Experiments are classified according to their input $Ro_{\text{cv,F}}$. Two output Rossby numbers, $Ro_{\omega}$ and $Ro_{\text{bulk}}$, are also included in the table. These are defined in Equation 9, and are intended to indicate the degree of rotational constraint at small and large scales, respectively. $Nu$ is the Nusselt Number (Equation 11), and $\Delta T$ is the radial temperature contrast. Both are evaluated for the convection zone (CZ; between $r_{\text{i}}=4$ and $r_{\text{o}}=5$). ‘MLT’ and ‘RMLT’ refer to ‘mixing length theory’ and ‘rotating mixing length theory’, respectively. The scalings for $Nu$ associated with these theories are given by Equations 12 and 13. $Re$ is the Reynolds number (Equation 8), and $KE_{\text{fluc}}$ is the fluctuating kinetic energy. Both are computed using the r.m.s. fluctuating (azimuthal average removed) velocity within the CZ. ‘VAC’ and ‘CIA’ refer to ‘Viscous-Archimedean-Coriolis balance’ and ‘Coriolis-inertial-Archimedean balance’, respectively. The VAC, CIA, and ‘Ultimate’ scalings are given by Equations 1517. Finally, the differential rotation classifications ‘FS’, ‘FSF’, and ’SF’ correspond to configurations with: (i) a fast equator and slow pole (Solar-like), (ii) a fast equator, slow mid-latitudes, and ‘fast’ polar vortex, and (iii) a slow equator and fast pole (anti-Solar).

### v2

@@ 3 Results
Figure 3: Snapshots of the perturbation temperature for six experiments. The experiments shown in the top row are in the rotationally-uninfluenced regime, those in the middle row are in the rotationally-influenced regime, and those in the bottom row are in the rotationally-constrained regime. In each panel, the inner surface corresponds to $r=4.2$ and the outer surface corresponds to $r=4.8$.
@@ 3.1 Simulation results
Figure 2a shows the shell-averaged temperature difference $\Delta T$ between the inner and outer radii of the CZ. For $Ro_{\text{cv,F}}<1$, $\Delta T$ displays a strong dependence on $Ro_{\text{cv,F}}$, increasing with $\Omega_{0}$ as the rotational constraint acts to throttle convective heat transport. For each $Ro_{\text{cv,F}}$, the scatter between experiments with different $Ra_{\text{F}}$ is very small. In the rotationally-influenced regime, the dependence of $\Delta T$ on $Ro_{\text{cv,F}}$ ($\Delta T\propto\Omega_{0}^{\frac{4}{5}}\,{\sim}\,Ro_{\text{cv,F}}^{-\frac{4}{5}}$) is consistent with rotating mixing length theory (RMLT; see Section 3.2; 1979GApFD..12..139S; 2014ApJ...791...13B), which is a diffusion-free theory. For the cases categorised as rotationally-constrained with $Ro_{\text{cv,F}}<0.1$, $\Delta T$ depends more strongly on $Ro_{\text{cv,F}}$ than predicted by RMLT, while remaining independent of the diffusivity (for the values of $Ro_{\text{cv,F}}$ for which there are multiple experiments). In the rotationally-uninfluenced regime where $Ro_{\text{cv,F}}>1$, the rotational constraint is lost, causing $\Delta T$ to become independent of $Ro_{\text{cv,F}}$. Our simulations yield results that are consistent with this expectation, and thus by extension are consistent with diffusion-free behaviour.
@@
The azimuthal velocity averaged within $\pm 5^{\circ}$ of the equator and throughout the depth of the CZ, $U_{\text{eq}}$, is shown in Figure 2b. The dependence of $U_{\text{eq}}$ on $Ro_{\text{cv,F}}$ we obtain is consistent with previous work (2014MNRAS.438L..76G). In particular, the transition between prograde and retrograde equatorial flow occurs when $Ro_{\text{cv,F}}\approx 0.1$, which corresponds to $Ro_{\omega}\approx 1$ (and $Ro_{l}\approx 1$; not shown). In the absence of a rotational constraint ($Ro_{\text{cv,F}}>1$), differential rotation does not develop ($U_{\text{eq}}\approx 0$). In the rotationally-influenced regime ($Ro_{\text{cv,F}}<1$; $Ro_{\omega}\gtrsim 1$), anti-Solar (retrograde) differential rotation develops, with an amplitude that increases as the rotation rate is increased. In this regime, the amplitude of the equatorial zonal flow is independent of the diffusivity. When rotation becomes dominant ($Ro_{\text{cv,F}}<0.1$; $Ro_{\omega}\lesssim 1$), the differential rotation switches direction from retrograde to prograde (Solar-like). In this regime, the data in general do not collapse onto a single curve for $U_{\text{eq}}$ vs. $Ro_{\text{cv,F}}$ indicating that, at least for the supercriticalities considered in the present work, the strength of $U_{\text{eq}}$ depends on the diffusivities. We note that there are only two values of $Ro_{\text{cv,F}}<0.1$ for which we ran a significant number of experiments with varying $Ra_{\text{F}}$ (5 experiments each for $Ro_{\text{cv,F}}=3.11\times 10^{-2}$ and $6.70\times 10^{-2}$). For the series with $Ro_{\text{cv,F}}=6.70\times 10^{-2}$, $U_{\text{eq}}$ does not vary significantly as the diffusivity is varied, but for the series with $Ro_{\text{cv,F}}=3.11\times 10^{-2}$ there is a diffusivity dependence.
@@
Figures 2c and d show the kinetic energy (KE) averaged over the volume of the CZ. Figure 2c shows $\text{KE}_{\text{fluc}}$, computed using the r.m.s. fluctuating velocity, and Figure 2d shows the ratio DRKE/KE, where KE is the total kinetic energy, and DRKE is the ‘differential rotation kinetic energy’, computed using the axially-symmetric component of the azimuthal velocity. The behaviour of $\text{KE}_{\text{fluc}}$ in Figure 2c tells a similar story to $U_{\text{eq}}$, namely that the kinematic properties of the flow display a weak dependence on the diffusivity for $Ro_{\text{cv,F}}>0.1$ (rotationally-uninfluenced or rotationally-influenced), but a stronger dependence once the influence of rotation becomes dominant. There are two additional features of Figure 2c,d that are worth emphasising. First, we note that in the rotationally-influenced regime, the dependence of $\text{KE}_{\text{fluc}}$ on $Ro_{\text{cv,F}}$ is consistent with the diffusion-free ‘Coriolis-inertial-Archimedean’ (CIA) scaling (e.g., 2016JFM...808..690G; see Section 3.2). By contrast, in the rotationally-constrained regime, $\text{KE}_{\text{fluc}}$ is diffusivity-dependent. However, the ratio DRKE/KE remains roughly diffusion-free. This is in large part due to the fact that the DRKE dominates the total KE in this regime, so that their ratio saturates at $\text{DRKE}/\text{KE}\approx 1$, although we note there is some compensation between the diffusivity dependence of the DRKE and $\text{KE}_{\text{fluc}}$ (i.e., for a given $Ro_{\text{cv,F}}$ an experiment with a larger $\text{KE}_{\text{fluc}}$ tends to have a larger DRKE).
@@
In the rotationally-uninfluenced regime, convective plumes extend outwards in the radial direction at all latitudes and the effect of rotation is undetectable. In this regime, the temperature field is organised into large ‘warm’ and ‘cool’ patches that are dominated by spherical harmonic degree $\ell=1$ or $2$. The morphology in the rotationally-influenced cases with $Ro_{\text{cv,F}}=3.11\times 10^{-1}$ is similar, although the convection at higher latitudes now exhibits a weak alignment with the rotation axis and some low-latitude features are elongated in the azimuthal direction. However, the impact of rotation is hard to detect, consistent with $Ro_{\omega}>1$. Figure 4 shows an additional case in the rotationally-influenced regime with $Ro_{\text{cv,F}}=1.44\times 10^{-1}$ (i.e., a slightly faster rotation rate), which more clearly shows that the convection is aligned cylindrically, and zonally-elongated at lower latitudes. We highlight that the pair of experiments with $Ro_{\text{cv,F}}=3.11\times 10^{-1}$ are morphologically very similar, as are the pair with $Ro_{\text{cv,F}}=1.44$, which is consistent with the inference from Figure 2 that the diffusivity has a limited influence on the dynamics in the rotationally-uninfluenced and rotationally-influenced regimes.
The effect of rotation in the rotationally-constrained regime is clear. In mid-latitudes and the polar regions, convection is aligned cylindrically (i.e., with the rotation axis). Closer to the equator, large, modulated thermal Rossby wave-like features can be identified, that propagate in the prograde direction (with respect to the mean zonal flow). These waves are present in all experiments with $Ro_{\text{cv,F}}<0.1$ but in none with $Ro_{\text{cv,F}}\geq 1$. The transition to a regime of equatorial dynamics dominated by wave-like features occurs when small and intermediate spatial scales become rotationally constrained ($Ro_{\omega}<1$). For the relatively turbulent cases shown in Figure 3 there is a strong polar vortex. This is not present in simulations that are more strongly dissipative (an example is given by the lower panel of Figure 4).
@@
Each of the scatter charts in this article use the marker shape to illustrate the occurrence of the SF, FSF, and FS configurations of differential rotation in parameter space (circles for SF, pentagons for FSF, and squares for FS). From Figure 1a–c, we identify that solar-like differential rotation (FS) is only obtained for the most diffusive experiments ($Ra_{\text{F}}\lesssim 10^{6}$), while more turbulent simulations with a strong rotational constraint exhibit FSF differential rotation. Figure 1d shows that the transition from a fast equator (either FS or FSF) to a slow equator occurs at the boundary between the rotationally-constrained and rotationally-influenced regimes, when $Ro_{\text{cv,F}}=0.1$, which corresponds to $Ro_{\omega}\approx 1$, consistent with previous work (2014MNRAS.438L..76G). The fact that all of the FS or FSF cases have $Ro_{\text{cv,F}}<0.1$, and all of the SF cases have $Ro_{\text{cv,F}}\geq 0.1$, implies that the transition from a fast equator to a slow equator is diffusion-free (even if the zonal flow structure in the fast-equator cases themselves displays a dependence on the diffusivity). We believe that this transition is associated with the existence of thermal Rossby wave-like features at low latitudes, which are present in all simulations in the rotationally-constrained regime but absent otherwise. However, a complete analysis of the zonal momentum budget is left for future work.
@@
Figure 6: Scaling results for NuNu (panels a and b) and ReRe (panels c and d) obtained with our experiments. In panels a) and b), grey and black dashed lines correspond to the mixing length theory (MLT) and rotating mixing length theory (RMLT) scaling laws given by Equations 12 and 13, respectively, and the black solid lines connect selected series of experiments with constant $Ro_{\text{cv,F}}$ (see caption of Figure 1 a–c for further information). In panel c), the scaling for ReRe predicted by VAC balance (Equation 17) is shown as a black dashed line. Finally, in panel d), the black dashed line shows the inertial scaling of rotating convection (associated with CIA balance; given by Equation 16), and the grey dashed line shows the diffusion-free, non-rotating ‘ultimate’ scaling $Re\propto Ra^{1/2}$ (Equation 15).
@@ 3.2.1 Nusselt number scaling
for $Pr=1$. A central assumption in MLT is that the total dimensional heat transport is independent of the diffusivities, and thus the scaling above is diffusion-free. We note that a simple dimensional analysis for a non-rotating system yields the scaling above under this assumption [^2] A diffusion-free scaling for a non-dimensional number (e.g., $Nu$ or $Re$) does not mean that the scaling parameter (e.g., $Ra$) is diffusivity-independent. Instead, it is a scaling that is consistent with the dimensional quantities (e.g., the dimensional radial temperature gradient for $Nu$) being diffusivity-independent. For the specific case of the MLT $Nu$ scaling, $Nu$ includes a factor of $\kappa^{-1}$ if $T$ is re-dimensionalised, and so the right-hand-side must also contain a factor of $\kappa^{-1}$, which implies a scaling $Nu\propto Ra^{\frac{1}{2}}Pr^{\frac{1}{2}}$ (which is $Ra^{\frac{1}{2}}$ for $Pr=1$).. For a rotating system, the analogous theory to MLT is rotating mixing length theory (RMLT; 1979GApFD..12..139S), which assumes that convection is dominated by modes with a growth rate well described by linear theory and amplitudes determined by a balance between their growth rate and the non-linear cascade rate. RMLT predicts that $Nu$ scales as
@@
As with Equation 12, this is a diffusion-free scaling. This scaling can also be obtained by combining the ‘diffusion-free heat transport assumption’ with a further assumption that the heat transport depends only on the supercriticality of the system $Ra/Ra_{\text{c}}$ (recall that $Ra_{\text{c}}\,{\sim}\,Ta^{\frac{3}{2}}$ for large $Ta$) (2012PhRvL.109y4503J; 2014ApJ...791...13B). We note that $Ra\approx Ra_{\text{F}}/Nu$, which yields the scaling $Nu\propto\left(Ra_{\text{F}}Ta^{-2/3}\right)^{3/5}Pr^{-1/5}$. Using the definition of $Ro_{\text{cv,F}}$, this can be re-written $Nu\propto Ra_{\text{F}}^{1/3}Pr^{1/3}Ro_{\text{cv,F}}^{4/5}$. Therefore, with $Pr$ constant, RMLT implies
@@
(i.e., as in MLT). Previous studies of boundary-driven Boussinesq convection in a spherical shell have not yielded MLT-like scaling for the Nusselt number in the regime where rotation is unimportant. Instead, shallower scaling exponents were obtained, due to the throttling of convection by the conductive heat transport in the boundary layers (e.g., 2015JFM...778..721G). Some studies have obtained RMLT-like scaling in the rapidly rotating regime, but only for extreme parameter values (e.g., 2016JFM...808..690G, wherein experiments with $Ta\gtrsim 10^{12}$ and $RaTa^{-2/3}\simeq 10$ exhibit a Nusselt number scaling in agreement with RMLT).
@@
Inspection of Figure 6a reveals that our experiments fall into three groups distinguished by different $Nu$ vs. $RaTa^{-2/3}$ scaling behaviours. When $Ro_{\text{cv,F}}\geq 1$ (indicated by a green marker colour), the Nusselt number follows the MLT scaling prediction (Equation 12). These experiments are those that we have categorised as rotationally-uninfluenced. For, $0.1\leq Ro_{\text{cv,F}}<1$ (indicated by blue marker colours), i.e., experiments in the rotationally-influenced regime, a steeper scaling is obtained that is consistent with RMLT (Equation 13). The transition between the MLT and RMLT scalings occurs when $Ro_{\text{cv,F}}=1$, which roughly corresponds to the point where $Ro_{\text{bulk}}=1$. The agreement between the scaling obtained by the rotationally-influenced experiments and RMLT is striking, given that the effect of rotation on the convective morphology in this regime can be relatively weak (e.g., Figure 3). These results (as well as those presented in the next subsection) inform our definition of the rotationally-uninfluenced and rotationally-influenced regimes. Finally, in the rotationally-constrained regime for which $Ro_{\text{cv,F}}<0.1$ (orange marker colours), the Nusselt number scaling departs from the RMLT branch occupied by the rotationally-influenced cases. This departure from RMLT is manifest as an offset from the RMLT branch (reduced $Nu$ for a given $RaTa^{-2/3}$) that increases as $Ro_{\text{cv,F}}$ is reduced. In Figure 6b, it is apparent that for constant $Ro_{\text{cv,F}}$, experiments in this regime roughly follow the $Nu\propto Ra^{1/2}$ scaling noted above (Equation 14; for simulations with sufficient supercriticality, $RaTa^{-2/3}\gtrsim 5$). This indicates that the convective heat transport is still independent of the diffusivity (consistent with the pseudodimensional $\Delta T$ shown in Figure 2a), and that the offset from the rotationally-influenced RMLT branch captures a dependence of the convective heat transport on $Ro_{\text{cv,F}}$ (a diffusion-free parameter) that is not accounted for by RMLT. Hereafter, we refer to this scaling as ‘offset RMLT’. As noted previously, the transition from the rotationally-influenced regime to this new regime occurs when $Ro_{\omega}\lesssim 1$, which indicates that all spatial scales are now affected by rotation, motivating our choice to refer to this regime as rotationally-constrained.
@@
When the same data are plotted against $Ra$ (Figure 6b), the MLT scaling obtained by the rotationally-uninfluenced regime can be identified for experiments with $Ta=10^{5}$, which unambiguously follow the $Ra^{1/2}$ prediction. The $Ta=10^{6}$ series follows the RMLT scaling for moderate supercriticality, before transitioning towards the MLT scaling at higher supercriticality, and the experiments at $Ta=10^{7}$ are approaching the RMLT scaling. Finally, the offset RMLT scaling identified for the rotationally-constrained regime is manifest as a steepening of the $Nu(Ra)$ relation for fixed $Ta$ (but, again, note that series of experiments with constant $Ro_{\text{cv,F}}$, joined by black solid lines, follow the scaling predicted by RMLT, in the form given by Equation 14).
@@
A preliminary analysis (not shown) of the latitudinal dependence of the heat budget and Nusselt number (cf. 2023JFM...954R...1G) suggests that the transition to the offset RMLT scaling that characterises the rotationally-constrained regime may be related to the change in the dynamics from a flow dominated by turbulent convection to a wave-dominated flow (we find wave-like features are always present when $Ro_{\text{cv,F}}<0.1$, but are absent when $Ro_{\text{cv,F}}\geq 0.1$). However, the offset scaling could alternatively arise due to interaction between the convection and the zonal flow, although this would have to be a ‘signed’ effect, since strong (anti-Solar) zonal flows are also present in the rotationally-influenced regime for which the standard RMLT scaling is recovered. We will explore this further in future work. We note that the offset scaling persists even for the rotationally-constrained experiments with the highest supercriticality. As a result, we do not believe that it arises because the rotationally-constrained simulations generally have lower supercriticality than those categorised as rotationally-influenced.
@@
Note. — Experiments are classified according to their input $Ro_{\text{cv,F}}$. Two output Rossby numbers, $Ro_{\omega}$ and $Ro_{\text{bulk}}$, are also included in the table. These are defined in Equation 9, and are intended to indicate the degree of rotational constraint at small scales, and the largest spatial scale, respectively. $Nu$ is the Nusselt Number (Equation 11), and $\Delta T$ is the radial temperature contrast. Both are evaluated for the convection zone (CZ; between $r_{\text{i}}=4$ and $r_{\text{o}}=5$). ‘MLT’ and ‘RMLT’ refer to ‘mixing length theory’ and ‘rotating mixing length theory’, respectively. The scalings for $Nu$ associated with these theories are given by Equations 12 and 13. $Re$ is the Reynolds number (Equation 8), and $\text{KE}_{\text{fluc}}$ is the fluctuating kinetic energy. Both are computed using the r.m.s. fluctuating (azimuthal average removed) velocity within the CZ. ‘VAC’ and ‘CIA’ refer to ‘Viscous-Archimedean-Coriolis balance’ and ‘Coriolis-inertial-Archimedean balance’, respectively. The VAC, CIA, and ‘Ultimate’ scalings are given by Equations 1517. Finally, the differential rotation classifications ‘FS’, ‘FSF’, and ’SF’ correspond to configurations with: (i) a fast equator and slow pole (Solar-like), (ii) a fast equator, slow mid-latitudes, and ‘fast’ polar vortex, and (iii) a slow equator and fast pole (anti-Solar).